### v1

# Revisiting Static Feature-Based Android Malware Detection

Abstract
The increasing reliance on machine learning (ML) in computer security, particularly for malware classification, has driven significant advancements. However, the replicability and reproducibility of these results are often overlooked, leading to challenges in verifying research findings. This paper highlights critical pitfalls that undermine the validity of ML research in Android malware detection, focusing on dataset and methodological issues. We comprehensively analyze Android malware detection using two datasets and assess offline and continual learning settings with six widely used ML models. Our study reveals that when properly tuned, simpler baseline methods can often outperform more complex models. To address reproducibility challenges, we propose solutions for improving datasets and methodological practices, enabling fairer model comparisons. Additionally, we open-source our code to facilitate malware analysis, making it extensible for new models and datasets. Our paper aims to support future research in Android malware detection and other security domains, enhancing the reliability and reproducibility of published results.
1 Introduction
Machine learning (ML) has become indispensable in computer security and is extensively used in literature for malware detection [7, 40, 37]. These models learn patterns from data, adapt to new threats, and scale far beyond the capabilities of traditional signature-based systems. However, despite their potential, the design and deployment of ML systems in security applications face significant challenges [6]. Key issues, as observed in prior works [6, 21, 51], include sampling bias, incorrect labels, data snooping, inappropriate threat models, and unsuitable baselines. Additionally, replication studies of these methods are often neglected, leading to claims that cannot be reliably reproduced. This lack of rigor in the design and evaluation process undermines the validity of many reported findings due to poor reproducibility [41]. Addressing these challenges requires a focus on thorough evaluation practices, ensuring that results are consistently reproducible across different environments and datasets.
Figure 1: Comparison of best and worst performing neural networks in Android malware detection using state-of-the-art active learning method from [15]. Models were initialized with 5 different random seeds. The average F1-score over months differs by 10.6%, and the False Negative Rate (FNR) differs by 13.9% between the two models, despite using the same hyperparameters. This highlights reproducibility challenges in machine learning research for Android malware detection.
An example of such reproducibility issues is illustrated in Figure 1. Here, we analyze the reproducibility of the results reported in [15] for Android malware detection using continuous learning. We replicated the experiments using the publicly available source code and dataset [25]. By using the provided best set of hyperparameters for the Drebin-based dataset [7], we conducted the experiments with five different random seeds. We then compared the best and worst-performing models on the test set. The best model achieved an average F1 score of 79.26% over the test months, while the worst model scored 68.65%, indicating a performance difference of 10.6% solely due to variations in random seed initialization. This example highlights how reliably reproducing prior results in Android malware detection can be fraught with challenges, which, if not adequately addressed, can hinder progress in this research field. Furthermore, such issues can lead to disparities between model performance reported in controlled environments and real-world applications, where the stakes are much higher [6].
In this paper, we explore the specific challenges associated with reproducibility and replicability in Android malware analysis, a critical area where these issues are particularly pronounced. By examining key aspects of the dataset and methodology, we shed light on the factors that undermine the credibility of research outcomes in this domain. We identify issues such as data duplicates, inadequate model selection, poor hyperparameter tuning, and flawed evaluation strategies. These contribute to disparities in results obtained under different experimental settings, leading to variability in model performance and thus hurting reproducibility. We systematically discuss how these factors cause irreproducibility and propose ways to mitigate them.
Our study primarily focuses on reproducing the results of a state-of-the-art method for active learning in Android malware detection [15]. However, we expand the scope by considering more diverse experimental settings and addressing general reproducibility issues in Android malware detection. We utilize six different models: Random Forest [13], Support Vector Machine [7], eXtreme Gradient Boosting [14], Multilayer Perceptron [45], Supervised Contrastive Classifier [55], and Hierarchical Contrastive Classifier [15], previously used in Android malware analysis. We analyze both offline learning, where the model is trained once, and active learning, where the model is periodically retrained with a subset of annotated samples. We perform detailed model selection and hyperparameter tuning for all models in both settings. Our analysis shows how data duplicates affect reproducibility, how adequate hyperparameter tuning of baseline models can yield surprising results compared to more complex models, and offers recommendations for enhancing reproducibility in future research. While we focus on Android malware detection, insights from this work can be extended to other domains, such as Windows malware detection [5, 54], and other security applications, like network intrusion detection [21, 41].
Key Contributions
  1. 1.
    We identify and discuss key issues and pitfalls in datasets and methodologies in Android malware analysis that lead to poor reproducibility of previously reported results. We propose solutions to address these pitfalls, aiming to improve reproducibility and enable fairer comparisons across different models.
  2. 2.
    We implement state-of-the-art machine learning methods for Android malware detection in both offline and active learning settings. Our rigorous study of model performance across diverse experimental settings reveals that simpler baseline methods can often outperform more complex models when properly tuned.
  3. 3.
    We open-source our code (to be made public upon publication) for the wider community to facilitate malware analysis. Our code is easily extensible for new models and tasks, supporting future research on Android malware detection and other security tasks to improve the reliability and reproducibility of published results.
2 Background & Related Work
Dataset Split Duration
Benign
Apps
Malicious
Apps
Total
Malware
Families
Feature
Dimension
Drebin Train 2019-01 to 2019-12 40,947 4,542 45,489 121 16,978
Validation 2020-01 to 2020-06 18,109 2,028 20,137 67
Test 2020-07 to 2021-12 30,797 3,631 34,428 71
APIGraph Train 2012-01 to 2012-12 27,472 3,061 30,533 104 1,159
Validation 2013-01 to 2013-06 21,310 2,366 23,676 115
Test 2013-07 to 2018-12 240,729 25,377 266,106 418
Table 1: Summary statistics of the datasets used in the study
2.1 Reproducibility & Replicability
We adhere to ACM’s definitions of reproducibility and replicability, standardized in 2020 [2]. Computational reproducibility refers to the ability of an independent team to achieve a study’s results using the original study’s artifacts [26]. This ensures that the reported findings are valid under specified conditions. Olszewski et al. [41] conducted a longitudinal study on the reproducibility of security papers published in Tier 1 venues (2013-2022), identifying challenges in reproducing results with author-provided artifacts and proposing mitigation strategies.
Replicability, in contrast, implies that an independent group can obtain the same results using artifacts they develop independently. Replicability studies generally involve different datasets and aim to confirm previous research results, considering the underlying system’s inherent uncertainty.
Our study examines both aspects of machine learning research for Android malware detection. For instance, the analysis shown in Figure 1 underscores reproducibility issues in prior research, even when author-provided artifacts are used, due to unaccounted randomness in published results. Additionally, we address dataset and methodological issues that must be resolved for a study’s claims to be replicable with independently developed artifacts. While previous works have analyzed ML design and datasets in security [6, 21, 51, 47], there is limited research on enhancing the replicability of ML results for malware detection.
In this paper, we use the term reproducibility to generally encompass both computational reproducibility and replicability.
2.2 Android Malware Detection
Android malware detection using machine learning classifies applications as benign or malicious [36]. Three main types of features are used: static, dynamic, and hybrid. Static features are obtained by analyzing the app’s source code or related information [7, 37, 58]. Specifically, the primary focus for Android applications is the APK file, the installation package, which includes components like AndroidManifest.xml and smali files obtained through decompilation. Dynamic features are acquired by observing the app’s behavior in real or emulated environments, such as sandboxes [29, 49, 57]. Hybrid features combine both static and dynamic characteristics [16, 38].
Extracted features are input into machine learning models like Random Forests [13], Support Vector Machines [7], and Multilayer Perceptrons [45] to classify apps as malware or benign. The choice of features and models depends on detection objectives and resources. Static analysis is fast and suitable for large-scale detection, whereas dynamic analysis can be more accurate but resource-intensive [3]. This work focuses on static-feature-based machine learning methods due to their prevalence in the literature [30, 42, 58, 8, 15]. However, the methodological flaws identified and recommendations apply to ML using other feature types as well.
Concept Drift:
Malware detection systems encounter challenges due to the evolving nature of malware, leading to concept drift. This drift can arise from new malware families, behavioral changes, evasion attempts, or updates in API semantics. Traditional supervised learning models trained on static datasets may become outdated as new malware types emerge, reducing detection accuracy [55, 17, 42, 53]. These models often fail to recognize emerging variants, resulting in higher false negatives. To address this, online or continuous learning methods can enable models to adapt to new data continuously and stay current with malware trends [30].
Reproducibility of Research in Android Malware Detection Using ML:
Previous studies have assessed issues in reported results for Android malware detection [59, 28], attempted to replicate prior research [18], or addressed limitations in practical environments [22]. However, these studies often rely on single global metrics like AUC or F1-score, lacking nuanced analysis of monthly performance variations, factors influencing reproducibility, and inherent variances in machine learning models due to their stochastic nature. Our work aims to address these gaps by thoroughly analyzing conditions that can lead to poor reproducibility in different experimental settings and how to mitigate them.
3 Study settings
3.1 Machine Learning
In this study, we consider two distinct machine-learning settings:
  1. 1.
    Offline Learning: In this classical supervised setting, the model is trained once on the training dataset and evaluated on the test set. Offline learning for malware detection involves using annotated data from a specific period (e.g., one year) to train a model and evaluate its performance on future unseen samples.
  2. 2.
    Active Learning: While offline settings may suffice when the data distribution remains constant, malware detection is susceptible to concept drift. Active learning addresses concept drift by periodically retraining the model [30, 58, 15]. This cost-effective approach focuses on annotating only the samples where the model is least confident, saving annotation resources. Although determining when to retrain remains challenging, a common practice is to retrain the model monthly. This provides a standard evaluation setting for active learning methods independent of drift detection techniques. We adopt this approach in our study.
We treat these settings separately because each requires different model selection methods, such as hyperparameter tuning, as noted in prior work [15]. We aim to analyze the factors influencing replicability in these settings and enhance the reliability of Android malware detection models.
3.2 Datasets
We utilize the publicly available dataset from [15], which includes two datasets based on static feature types: Drebin [7] and APIGraph [58]. Hereafter, we refer to these as the Drebin [^1] This dataset is collected from and referred to as ”Androzoo”[1] in  [15] but since it uses the same feature sets as original Drebin dataset [7], we refer to it as Drebin. and APIGraph datasets. Drebin uses eight different sets of features, including access to hardware components, requested permissions, app components, intents, usage of restricted API calls, used permissions, suspicious API calls, and network addresses. APIGraph employs API semantics to cluster similar APIs, significantly reducing the feature space. The resulting feature set is binary for both datasets, indicating whether a feature is present or absent for a specific sample.
The Drebin dataset contains applications collected from 2019 to 2021, while the APIGraph dataset includes applications from 2012 to 2018. Both datasets address temporal and spatial biases present in malware datasets [6, 42]. They consist of approximately 90% benign and 10% malware applications collected from each month, reflecting their real-world ratios. Samples are ordered and approximately evenly distributed over their respective periods. Table 1 provides a summary of the datasets.
Following [15], we use data from the first year as the training set, the next six months for validation, and the remaining months for testing. For offline, only the training set data is used to train the model. In the active learning setup, this data is used to train the initial model, which is then periodically retrained using data from subsequent months from the validation and test set. The validation set is used to tune the model hyperparameters in both experimental settings.
(a) Drebin: Offline learning deduplication
(b) Drebin: Active learning deduplication
(c) APIGraph: Offline learning deduplication
(d) APIGraph: Active learning deduplication
Figure 2: Fraction of unique samples retained after the deduplication process
4 Pitfalls
In this section, we examine the challenges in the existing literature on malware detection using machine learning. We focus on dataset and methodological issues that impact the reproducibility of reported results.
4.1 Data Duplicates
Issues of duplicates in Android malware datasets have been reported in prior studies [28, 59]. The study by Irolla et al. [28] analyzed duplicates in the original Drebin dataset [7] using opcode sequences and found that 50.65% of malware samples were unique, with the rest being duplicates. They demonstrated that removing duplicates can alter the performance rankings of different machine learning models. Zhao et al. [59] examined duplicates in four malware datasets using APK, DEX, opcode sequences, and API calls to identify them. They also evaluated the impact of duplicates on multiple machine-learning models.
Our study differs from these works in two key ways. First, neither study provided a time-aware analysis of duplicates’ effects on malware detection models. Time-aware metrics are recommended for malware detection [42], and results based only on global metrics (such as F1-score) may not accurately reflect the situation, as the level of duplicates can vary monthly. We will demonstrate this variability later. Second, we evaluate the impact of duplicates in both offline and active learning settings, whereas previous studies focused only on offline learning. These settings require different methods to address duplicates, which we discuss later.
4.1.1 Defining Duplicates
We define two samples as duplicates if their feature sets are identical after the feature extraction process, meaning they are presented as identical inputs to the machine learning model. Specifically, two samples $X$ and $Z$ are considered duplicates if $x_{1}=z_{1},x_{2}=z_{2},\ldots,x_{n}=z_{n}$ for an $n$-dimensional feature vector:
$X=(x_{1},x_{2},\ldots,x_{n}),\quad Z=(z_{1},z_{2},\ldots,z_{n})$
This definition ensures that duplicate analysis is independent of how features are derived, unlike previous studies [28, 59]. It is important to note that this definition does not include "near-duplicates," such as samples from the same malware family that may have similar behavior but slightly different feature sets due to changes in benign code.
The presence of duplicate samples can bias machine learning models by causing specific training samples to be redundant and skewing performance metrics on the test set [4]. If a sample appears in both the training and test sets, it can artificially inflate test set performance, thus failing to reflect the model’s true generalization abilities accurately. Therefore, it is crucial to consider duplicates when evaluating machine learning models.
4.1.2 Deduplication
The process of deduplication ensures that the dataset contains unique samples. We define two types of deduplication based on different experimental settings in our study:
Offline Learning Deduplication: In offline learning, the model is trained once on a training set and evaluated on validation and test sets. Deduplication involves selecting unique samples within each data split. The model is evaluated monthly, requiring two levels of deduplication. First, we remove samples that appear in earlier splits. For the validation set, we remove duplicates from the training set. We remove duplicates from the training and validation sets for the test set. Within each split, we keep only the earliest occurrence of a sample. This ensures that all samples across the training, validation, and test sets are unique, retaining the earliest samples based on appearance time.
Active Learning Deduplication: When retraining the model monthly, the definition of duplicates changes due to continuous model updates, which can affect behavior, such as due to catastrophic forgetting [44]. Following [15], we assume that training data is always available. We remove duplicate samples from the training split, which is similar to offline learning. For the validation and test sets, duplicates are removed only within the same month, not across different months. This approach allows duplicates to appear in multiple months of the test set but ensures each month contains unique instances. This approach is essential for active learning as the model continuously evolves. If a sample appears in two months, the model should still recognize it accurately. By maintaining unique samples each month, we assess the model’s performance fairly and prevent duplicates from affecting retraining.
4.1.3 Duplicate Statistics in the Datasets
Figure 2 illustrates the proportion of unique samples in the datasets after the deduplication process in both experimental settings. Ideally, this fraction would be 1 for both malware and benign apps, indicating no duplicates. However, our analysis shows that many duplicates are present, with varying degrees across different splits. Notably, around 50% of the training set for the Drebin and APIGraph datasets consists of duplicate samples in benign and malware categories.
In the validation and test sets, we observe two key patterns regarding duplicates. First, the fraction of duplicate samples varies monthly. Second, malware samples tend to contain more duplicates than benign samples, with certain months particularly affected. For instance, in the first test month of the Drebin dataset (July 2020), only 6.30% of malware samples are unique in the offline-learning setting, and 9.96% are unique in the active-learning setting. This significantly changes the benign-to-malware sample ratio in the deduplicated dataset. Initially, the ratio for this month was 8.81, but it increased to 62.45 for offline learning and 50.65 for active learning deduplication.
We demonstrate in Sections 6 and 7 how the presence of duplicates can undermine the replicability of reported results in Android malware detection, leading to biased evaluations of model performance.
4.2 Model Selection Strategy
A common approach for evaluating machine learning models involves splitting the dataset into training, validation, and test sets. The validation set is essential for model selection and hyperparameter tuning to prevent biased parameter choices [6]. However, previous research on Android malware detection often omits a separate validation set, leading to incomplete evaluations [6]. Our study adopts the same train, validation, and test split as in [15] for offline and active learning settings.
Another common issue is the inadequate tuning of baseline methods [6]. This often results in insufficient hyperparameter tuning for simpler models compared to more complex ones. For example, prior research has shown that appropriately tuned tree-based methods frequently outperform deep neural networks on tabular datasets [24, 39], even though neural networks excel in tasks like computer vision and natural language processing [33, 27, 50, 20].
In Android malware detection, various models have been used in literature, such as support vector machines (SVM), random forests, gradient-boosted decision trees, and deep neural networks [7, 37, 42, 58, 15]. However, these models might not have been adequately calibrated due to the absence of a validation set or insufficient hyperparameter searches. Our analysis focuses on extensive hyperparameter searches for different baseline methods, maintaining the same hyperparameter budget for different models to enable a fair comparison between the models. We detail our procedure in Section 5.3 and show how it can lead to significantly different conclusions about model performance in Section 6.4.
4.3 Accounting for Variance in Performance
Some machine learning models, particularly neural networks, introduce additional stochasticity independent of their hyperparameters. They are sensitive to initial random weight initialization, leading to variations in results based on the training system rather than model parameters [11, 43]. To address this, the deep learning community averages performance across multiple random seeds on the test set and reports the mean performance [12]. This approach captures performance variance, providing a more robust estimate of the model’s capabilities and ensuring reported metrics are not artifacts of a single training instance. However, previous studies on Android malware detection often overlook this variation, resulting in irreproducible outcomes when experiments use different random seeds. We discuss this in detail in Section 6.2.
4.4 Delayed Evaluation of Models
Splitting the dataset into non-overlapping temporal train, validation, and test sets aids in better model selection. However, this approach often results in the test dataset being delayed from the training dataset. While this is not an issue for classical supervised learning, which assumes an independent and identically distributed setting, it poses challenges for malware detection. The test dataset can be significantly separated from the training dataset by the duration of the validation set, potentially causing concept drift during this period.
For instance, in our study and in [15], there is a six-month gap between the training and test datasets. Using the model trained solely on the training data to evaluate the test set directly might not reflect the practical deployment scenario. Ideally, the model should be trained with more recent data before deployment [6].
To address this characteristic of malware datasets, we propose two different evaluation settings:
  1. 1.
    Merged Training: In this setting, we merge the training and validation datasets after hyperparameter tuning and evaluate the model on the test dataset. We propose that this setting provides a fairer estimate of model performance for offline learning, where the model is already delayed when the test month starts. Note that we do not employ a separate validation set during the merged training phase.
  2. 2.
    Holdout Training: This is the standard setting where the validation dataset is not merged with the training dataset. The test set is separated from the training dataset in this setting, and performance may be underestimated due to concept drift.
We show the effect of these two evaluation settings in Sections 6 and 7.
5 Experiments
5.1 Machine Learning Models
We utilize six different models commonly used in prior works on Android malware detection:
  1. 1.
    Random Forest (RF): Random Forest [13] is an ensemble learning method for classification that constructs multiple decision trees during training and outputs the mode of their classes. It has been used in Android malware detection [46, 37].
  2. 2.
    Support Vector Machine (SVM): SVM is a classification algorithm that finds the hyperplane that best separates classes in the feature space. It has been widely used for Android malware detection, particularly with Drebin features [7, 30].
  3. 3.
    eXtreme Gradient Boosting (XGBoost): XGBoost [14] is a scalable and efficient gradient boosting method that excels in performance on various tabular datasets [24], although it has not shown consistent success in Android malware detection [15].
  4. 4.
    Multilayer Perceptron (MLP): An MLP is a type of neural network composed of multiple layers of neurons, with each layer fully connected to the next [45]. MLPs are used for a variety of tasks, including classification and regression, and have been employed in several prior works for Android malware detection [42, 56, 32].
  5. 5.
    Supervised Contrastive Classifier (SCC): Contrastive learning effectively detects and adapts to concept drift in Android malware detection [55, 15]. This approach extracts meaningful representations by distinguishing between similar and dissimilar instance pairs. It assumes that similar instances should be closer in the embedding space while dissimilar ones should be further apart [52, 31].
    We implement a method based on triplet learning, focusing on the main classes—benign and malware—with a triplet loss. In our approach, a sample is considered positive if it belongs to the same class as the anchor and negative if it belongs to a different class. Unlike previous works that treat each malware family as a distinct class, leading to $m+1$ classes where $m$ is the number of malware families, our method only considers the binary case. This is more annotation-friendly, as it does not require separate malware family annotation.
    The learning process combines a supervised binary cross-entropy loss and a triplet margin loss. The binary cross-entropy loss is defined as:
    $\mathcal{L}_{\text{bce}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$
    where $y_{i}$ is the true label and $\hat{y}_{i}$ is the predicted probability.
    The triplet margin loss is given by:
    $\mathcal{L}_{\text{triplet}}=\frac{1}{N}\sum_{i=1}^{N}\max\big{(}$ $0,\|\text{enc}(x_{i}^{a})-\text{enc}(x_{i}^{p})\|_{2}^{2}$
    $-\|\text{enc}(x_{i}^{a})-\text{enc}(x_{i}^{n})\|_{2}^{2}+\text{m}\big{)}$
    where $x_{i}^{a},x_{i}^{p},x_{i}^{n}$ are the anchor, positive, and negative samples respectively, $\text{enc}(\cdot)$ is the encoder function, and $m$ is the margin parameter.
    The combined loss function is:
    $\mathcal{L}=\lambda\cdot\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{triplet}}$
    where $\lambda$ is a weight parameter balancing the two loss components.
  6. 6.
    Hierarchical Contrastive Classifier (HCC): Chen et al. [15] introduced an enhanced supervised contrastive classifier featuring a hierarchical contrastive loss function. The core idea is to ensure that samples from the same malware family are more similar than those from different families in the feature embedding space. Using an encoder network enc, the embeddings of two benign or two malicious samples from different families should satisfy $\|\text{enc}(x_{1})-\text{enc}(x_{2})\|_{2}\leq m$. For samples from the same malware family, the embeddings should be even more tightly constrained, while embeddings of one malicious and one benign sample should be highly dissimilar, $\|\text{enc}(x_{1})-\text{enc}(x_{2})\|_{2}\geq 2m$.
    Let $d_{ij}$ denote the Euclidean distance between two samples $i$ and $j$ in the embedding space, defined as $d_{ij}=\|\text{enc}(x_{i})-\text{enc}(x_{j})\|_{2}$. The hierarchical contrastive loss is:
    $\mathcal{L}_{hc}(i)=$ $\frac{1}{|P(i,y_{i},y_{i}^{\prime})|}\sum_{j\in P(i,y_{i},y_{i}^{\prime})}\max(0,d_{ij}-m)$
    $+\frac{1}{|P_{z}(i,y_{i},y_{i}^{\prime})|}\sum_{j\in P_{z}(i,y_{i},y_{i}^{\prime})}d_{ij}$
    $+\frac{1}{|N(i,y_{i})|}\sum_{j\in N(i,y_{i})}\max(0,2m-d_{ij})$
    The first term ensures positive pairs, such as (benign, benign) or (malicious, malicious), are close but not overly constrained, penalizing distances only if they exceed $m$. The second term enforces similarity within the same malware family. The last term aims to separate benign and malicious samples by at least $2m$.
    Like the SCC, this loss is combined with binary cross-entropy loss to train the model. Note that this is the only model that requires access to malware family labels to train the model.
5.2 Active Learning Sample Selection Strategy
Active learning focuses on selectively choosing the most informative samples from a large pool of unlabeled data to label, aiming to improve model performance with minimal labeled data. We employ the baseline uncertainty sampling strategy [34] for all models except the HCC. The core idea of uncertainty sampling is to select samples about which the model is least confident. For our models, this involves computing $1-\text{probability}$ (e.g., the softmax output for neural networks) for binary classes and selecting samples with the highest uncertainty.
For HCC, we use pseudo-label sample selection as described in [15]. This method uses model predictions, or pseudo labels, as ground truth to compute the loss function. It calculates the pseudo-loss for contrastive learning by comparing embeddings of test samples against nearby samples in the training dataset. The method then combines the contrastive loss with the binary cross-entropy loss, similar to the one used in training, to identify uncertain samples.
5.3 Hyperparameter Tuning
We conduct hyperparameter tuning independently for two experimental setups: offline learning and active learning.
In the offline learning setup, each model is trained once on the training dataset, using the validation set to select optimal hyperparameters. Since the model is trained only once, the sample selection strategy (Section 5.2) does not affect this setup. We use random search [10] for hyperparameter tuning, with a fixed budget of 200 searches per model on the validation set. Performance plateaued after 100 iterations for most models. This approach allows for a fair comparison across models and accounts for performance variations due to different hyperparameter search spaces. We perform hyperparameter tuning separately for both the duplicated and deduplicated datasets.
In the active learning setup, we follow the strategy from [15]. The validation data is divided into six-month intervals. We retrain the model using active learning for each interval and evaluate it in the subsequent month. Most hyperparameters, such as batch size for neural networks, remain consistent across the initial training and each retraining phase (a total of five, as no retraining is needed for the last validation month). However, parameters like the number of training epochs during retraining may vary. We use random search with a budget of 100 searches per model and a fixed annotation budget of 50 samples per month, similar to [15], for active learning hyperparameter tuning.
For both setups, we primarily use hyperparameters from previous studies [15], but we have expanded the search space for specific models, such as XGBoost, to improve performance.
Detailed hyperparameters and their influence on validation set performance are provided in the Appendix B.
5.4 Performance Evaluation
After hyperparameter tuning, we evaluate the models on the test set. The active learning setup includes retraining the model each test month and evaluating it the following month. We use the same performance metrics as in [15]: average F1-score, False Positive Rate (FPR), and False Negative Rate (FNR). These metrics differ slightly from the Area Under Time (AUT) metric in [42] but are highly correlated since the interval is consistently one month. Thus, conclusions from both metrics are similar.
Four models, including the random forest and the three neural networks (MLP, SCC, HCC), depend on random seeds for initialization. This reliance makes reproducibility challenging when reimplementing on different platforms, even with identical hyperparameters [60]. To mitigate this issue, we conduct experiments using optimal hyperparameters on the test set with five different seeds, reporting the mean and standard deviation of the metrics. This approach highlights the variability in reproducing model performance across different environments.
While XGBoost can introduce randomness through its sampling methods, we use default parameter values (i.e., 1) to ensure deterministic outcomes for the same model parameters [19]. We confirm deterministic results by running experiments on a different system with the same XGBoost library version.
5.5 Experimental Settings
We summarize the different experimental settings used in our analysis:
  1. 1.
    Learning Settings:
    • Offline learning: The model is trained once.
    • Active learning: The model is retrained monthly with a fixed annotation budget. For RF, SVM, and XGBoost, we train from scratch, while for neural networks, we continue fine-tuning the previous model, following recommendations from [15].
  2. 2.
    Duplicates:
    • Duplicated: We use the original datasets containing duplicated samples in the train, validation, and test splits.
    • Deduplicated: We deduplicate the datasets following the procedures in Section 4.1.2 before conducting the experiments.
  3. 3.
    Delayed Evaluation: Based on whether validation data is merged with training data after hyperparameter tuning(4.4):
    • Merged Training: Merge validation data with training data before evaluating the model on test data.
    • Holdout Training: Train the model using only training data.
    Both settings share the same set of hyperparameters.
  4. 4.
    Active Learning Budget: We evaluate two monthly annotation budgets for the active learning setup: 50 and 100 samples. Both settings use the model optimized with a 50-sample budget for hyperparameter tuning.
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 46.7±0.99 0.14±0.00 66.8±0.74 41.5±0.30 0.08±0.00 72.5±0.25 39.4±1.31 0.16±0.03 72.4±1.03 33.2±1.01 0.14±0.04 78.7±0.71
SVM 63.6±0.00 0.80±0.00 47.0±0.00 46.6±0.00 0.60±0.00 65.6±0.00 47.2±0.00 0.92±0.00 63.2±0.00 33.8±0.00 0.70±0.00 76.7±0.00
XGBoost 62.9±0.00 0.18±0.00 49.9±0.00 59.9±0.00 0.41±0.00 51.7±0.00 48.0±0.00 0.16±0.00 64.6±0.00 44.9±0.00 0.66±0.00 65.6±0.00
MLP 66.2±3.69 0.41±0.08 45.8±4.10 53.6±2.08 0.39±0.08 59.5±2.12 44.3±2.07 0.87±0.06 66.7±1.72 40.9±1.05 0.57±0.04 69.8±0.82
SCC 74.7±0.78 0.53±0.09 35.0±1.51 57.4±0.44 0.63±0.25 53.7±2.11 45.6±1.18 0.70±0.11 65.7±1.35 37.3±0.26 0.63±0.10 73.0±0.61
HCC 68.4±1.73 0.82±0.21 41.3±2.52 56.7±0.79 1.04±0.12 50.5±0.76 44.9±1.60 0.77±0.11 66.2±1.60 35.8±0.65 0.84±0.15 73.5±1.14
Table 2: Performance comparison of different models on Drebin Dataset for offline learning
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 64.6±1.02 0.23±0.00 49.6±1.08 61.3±0.92 0.50±0.03 51.7±0.79 45.3±5.2 0.12±0.02 68.5±4.50 46.8±3.89 0.34±0.08 67.0±3.76
SVM 76.0±0.00 1.20±0.00 30.7±0.00 69.7±0.00 1.41±0.00 34.9±0.00 74.5±0.00 0.86±0.00 34.7±0.00 66.3±0.00 1.42±0.00 39.5±0.00
XGBoost 74.1±0.00 1.45±0.00 31.7±0.00 68.8±0.00 1.60±0.00 34.7±0.00 70.6±0.00 1.07±0.00 38.5±0.00 66.5±0.00 0.97±0.00 42.7±0.00
MLP 68.6±2.28 0.85±0.26 41.9±3.77 63.0±2.73 0.96±0.37 46.32±4.9 55.6±4.26 0.90±0.20 56.2±4.85 57.0±3.52 1.67±0.64 49.9±6.78
SCC 66.1±2.79 1.06±0.26 43.8±3.94 66.2±2.09 2.11±0.57 35.2±1.98 70.36±2.9 1.79±0.16 35.3±3.33 57.0±3.52 1.67±0.64 49.9±6.78
HCC 73.3±2.37 1.35±0.15 33.3±3.90 66.0±1.75 0.92±0.17 42.8±2.66 64.2±5.17 1.60±0.20 43.8±6.61 62.6±2.74 1.59±0.26 43.8±4.13
Table 3: Performance comparison of different models on APIGraph Dataset for offline learning
6 Offline Learning Results & Analysis
6.1 Malware Detection Performance
We present the results of six models in Table 2 for the Drebin dataset and Table 3 for the APIGraph dataset. These results encompass the various experimental settings discussed previously. Key findings are summarized as follows:
  1. 1.
    Impact of Duplicates on Model Performance: Deduplicated datasets generally perform worse than those with duplicates, indicating that duplicates can inflate performance metrics. For instance, in the Drebin dataset’s merged training setting, the F1-score differences for the SVM and SCC models are 17.0 and 17.3, respectively, between duplicate and deduplicated datasets.
  2. 2.
    Benefits of Merged Training: Merging validation data with training data before final evaluation consistently enhances performance. In the Drebin deduplicated dataset, the average F1-score improvement ranges from 8.3 (RF) to 20.9 (HCC), demonstrating the benefits of additional training data and its alignment with practical applications.
  3. 3.
    Duplicates Influence Model Preference: Duplicates can skew model comparisons. For example, in the Drebin dataset merge-training setting, SCC outperforms HCC by 6.3 in average F1-score. However, after deduplication, XGBoost outperforms SCC by 2.5 in average F1-score.
  4. 4.
    XGBoost’s Optimal Performance: Adequately tuned, XGBoost achieves the highest average F1-score in 3 out of 4 deduplicated datasets and remains competitive in others, suggesting it is a strong baseline model. While SVM performs well on the APIGraph dataset, its performance declines notably on the Drebin dataset after deduplication.
  5. 5.
    Increased Variability Due to Duplicates: Models reliant on random initialization (RF, MLP, SCC, HCC) exhibit greater performance variance on duplicate datasets in most experimental settings. Duplicates not only inflate results but also increase variability, affecting reproducibility. For the APIGraph dataset in the duplicate holdout training setting, HCC’s average F1-score ranges from 57.7 to 70.7, showing a 13-point difference due to random seed variations. We discuss this issue further below.
Figure 3: F1-score over the test months on the APIGraph for (left) deduplicated (right) duplicated datasets
(a) F1-score on deduplicated data
(b) F1-score on duplicated dataset
(c) FPR on duplicated dataset
(d) FNR on duplicated dataset
Figure 4: Performance in test months on Drebin datasets for Merged Training setting
6.2 Duplicates and Variance in Performance
Figure 5: Frequency of Unique and top-10 Duplicate malware samples for the month 2015-11 on APIGraph dataset (left) and the month 2021-01 on Drebin dataset (right)
Figures 3 and 4 show the month-by-month performance of four models (SVM, XGBoost, MLP, HCC) on the APIGraph and Drebin datasets under the merged training setting for both duplicated and deduplicated datasets. Key conclusions from the results are:
  1. 1.
    Impact of Duplicates on Performance Trends: On both datasets, models generally exhibit a downward performance trend over time due to concept drift. However, this trend is less evident in the original datasets with duplicates, where sudden fluctuations occur. For example, all models experienced a performance drop in November 2015 on the APIGraph dataset, explained by duplicate inputs, as shown in Figure 5 (left). Of the 488 malware samples that month, 190 have identical input features. If a model fails to detect one correctly, it fails for all, resulting in a significant performance drop.
  2. 2.
    High Variance Due to Duplicates: As seen in Figure 4(b), duplicates can cause high variance in model performance. Figure 4(d) shows that this effect is primarily due to variance introduced by false negative rates (FNRs). Figure 5(d) illustrates the frequency of unique and top-10 duplicated malware samples for January 2021 on the Drebin dataset. Specifically, 348 out of 385 samples are duplicates of the same input feature, resulting in significant variance for models relying on random initialization, such as neural networks. Two of five MLPs trained with different seeds achieved high F1 scores (93.7 and 94.0), while the others scored much lower (21.4, 24.6, 24.8), indicating high fluctuations due to duplicates. A theoretical classifier correctly identifying all malware samples this month has an FNR of 0. If it fails to detect this single duplicated sample, the FNR rises to 90.4%. Conversely, a classifier detecting no samples has an FNR of 100%, but identifying only the duplicated sample reduces the FNR to 9.6%. This highlights the need to deduplicate datasets for models relying on random initialization to prevent performance disparities caused by "lucky" or "unlucky" initializations.
  3. 3.
    Importance of reporting performance with multiple seeds: Although deduplication reduces extreme variance in model performance for certain months, neural networks can still show performance variation due to random initialization, as seen in Figure 4(a). For instance, in four test months (2021-01, 2021-08, 2021-10, 2021-12), the MLP had a standard deviation of average F1-scores greater than 5, even on the deduplicated Drebin dataset. Therefore, accounting for this randomness when reporting neural network performance is essential.
6.3 Duplicates & Model Selection
Figure 6: Effect of balancing the weight of the classes during learning for XGBoost model on different datasets
The presence of duplicates can also affect hyperparameter tuning and model selection. Figure 6 illustrates this for the XGBoost model, where we adjust the "scale_pos_weight" parameter to address the class imbalance. The positive class weight (malware) is set to 1 in the imbalanced setting. For the balanced setting, it is adjusted to the ratio of benign to malware samples in the training dataset. Balancing the weight improves validation performance for the duplicated dataset in both cases. However, it may reduce performance on the deduplicated APIGraph dataset. This suggests duplicates can cause some hyperparameters to behave differently for the same model.
Figure 7: Comparison of feature weights for the top 50 features in XGBoost models trained on the deduplicated Drebin dataset and their corresponding weights in the model trained on the original duplicated dataset.
Figure 7 illustrates how duplicates can influence the model to weigh features differently during training. It compares feature importance between two XGBoost models trained on the deduplicated and the original duplicated Drebin dataset. We extracted the top 50 features based on XGBoost’s feature importance from the deduplicated dataset and evaluated their corresponding importance in the model trained using the duplicated dataset. The figure reveals a disparity in the distribution of feature importance between the two models.
Figure 8: Effect of number of boosting rounds during learning for XGBoost model on different datasets
6.4 Effective Model Selection for Robust Baselines
Our experiments with offline learning show that the XGBoost model achieves superior performance across various experimental settings, contrasting with previous results [15]. The key reason is the expanded hyperparameter search space. We included more hyperparameters and broadened the search range, as detailed in Appendix B. A critical change was the number of boosting rounds considered. Chen et al. [15] tested rounds ranging from 10 to 100. However, as shown in Figure 8, XGBoost often benefits from more boosting rounds, particularly on the APIGraph dataset, where 400 rounds yielded the best results for duplicated and deduplicated datasets. This underscores the importance of proper hyperparameter tuning to establish stronger baseline models [6].
Model Budget=50 Budget=100
Merged Training Holdout Training Merged Training Holdout Training
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 61.9±0.73 0.13±0.02 53.0±0.88 62.7±1.11 0.11±0.00 52.4±1.11 62.8±1.24 0.13±0.01 52.1±1.30 63.6±0.47 0.12±0.03 51.3±0.40
SVM 64.4±0.00 0.22±0.00 48.7±0.00 59.6±0.00 0.23±0.00 54.1±0.00 68.1±0.00 0.24±0.00 44.6±0.00 63.9±0.00 0.26±0.00 49.1±0.00
XGBoost 82.1±0.00 0.17±0.00 27.2±0.00 81.3±0.00 0.14±0.00 29.0±0.00 82.8±0.00 0.16±0.00 26.4±0.00 81.6±0.00 0.17±0.00 28.0±0.00
MLP 67.9±0.65 0.49±0.03 43.3±0.70 61.8±0.54 0.35±0.02 51.0±0.49 70.2±0.42 0.38±0.03 40.6±0.46 65.7±3.37 0.36±0.06 46.2±4.17
SCC 71.0±0.64 0.59±0.04 37.4±0.29 68.2±1.83 0.50±0.07 41.6±2.21 73.7±1.20 0.57±0.10 33.7±1.35 72.1±0.88 0.48±0.07 37.0±1.44
HCC 73.5±0.27 0.66±0.04 32.8±0.85 68.5±3.38 0.57±0.05 40.3±4.37 75.6±0.96 0.68±0.06 29.4±1.07 72.6±2.10 0.57±0.02 35.2±2.64
Table 4: Performance of different models on Drebin Dataset for active learning with 50 & 100 annotation budget per month
Model Budget=50 Budget=100
Merged Training Holdout Training Merged Training Holdout Training
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 87.1±0.15 0.27±0.00 19.4±0.28 87.4±0.20 0.30±0.01 18.6±0.34 89.0±0.11 0.26±0.01 16.4±0.13 89.4±0.11 0.28±0.01 15.6±0.11
SVM 83.7±0.00 0.60±0.00 21.4±0.00 84.4±0.00 0.67±0.00 19.2±0.00 84.8±0.00 0.63±0.00 19.0±0.00 85.2±0.00 0.75±0.00 16.9±0.00
XGBoost 88.4±0.00 0.36±0.00 16.3±0.00 88.7±0.00 0.36±0.00 15.8±0.00 90.2±0.00 0.33±0.00 13.7±0.00 90.4±0.00 0.32±0.00 13.4±0.00
MLP 85.7±0.67 0.47±0.03 19.6±1.12 85.7±0.40 0.47±0.05 19.5±1.10 87.4±0.29 0.47±0.02 16.6±0.49 87.2±0.09 0.49±0.04 16.6±0.60
SCC 85.6±0.37 0.85±0.07 15.1±0.27 86.0±0.71 0.81±0.08 14.8±0.31 86.1±0.38 0.95±0.07 12.8±0.61 85.7±2.80 2.03±2.51 12.9±0.48
HCC 87.1±0.36 0.64±0.05 15.6±0.95 86.7±0.15 0.56±0.03 17.2±0.36 87.5±0.78 0.56±0.03 15.5±1.11 87.4±0.27 0.59±0.02 15.6±0.14
Table 5: Performance of different models on APIGraph Dataset for active learning with 50 & 100 annotation budget per month
7 Active Learning Results & Analysis
In the active learning setting, duplicates pose additional challenges beyond those in offline learning. Since active learning relies on human experts to select samples for annotation, increasing the diversity of selected samples is crucial [48]. Duplicates can lead to redundant selections, wasting the annotation budget. A simple heuristic can identify previously annotated samples and replace them with their existing annotations. Thus, using duplicated datasets in an active learning setting is impractical, and our analysis primarily focuses on deduplicated datasets.
7.1 Malware Detection Performance
We present the results of the six models in Tables 4 and 5 for the Drebin and APIGraph datasets, discussing merged and holdout training settings. Unlike offline learning, in active learning, the model is continuously updated during validation months using a subset of samples. Key findings are summarized below:
  1. 1.
    XGBoost’s Superior Performance: Similar to offline learning, XGBoost achieves the highest average F1 score across all experimental settings in active learning. This is especially evident in the challenging Drebin dataset, where it surpasses the second-best model by 8.1% in the merged training setting. This highlights the importance of selecting a strong baseline model and underscores XGBoost’s effectiveness in Android malware detection.
  2. 2.
    Benefits of Contrastive Learning for Neural Networks: The neural networks trained with contrastive learning, SCC, and HCC outperform baseline neural networks in all settings, indicating the effectiveness of contrastive learning in enhancing robustness to concept drift. HCC further outperforms SCC, demonstrating the advantages of hierarchical contrastive learning and the pseudo-loss sample selector introduced in [15]. However, more improvements are needed to match XGBoost’s performance.
  3. 3.
    Random Forest as a Competitive Baseline on APIGraph: Even simpler tree-based methods like Random Forest achieve competitive performance on the APIGraph dataset, often surpassing more complex neural networks. With consistently low false positives, RF is a strong candidate model for this dataset, illustrating that simpler models can compete effectively with complex ones in a compact feature space.
  4. 4.
    Performance Gains from Merged Training: Merging the validation set with the training set can enhance performance on test data, particularly for neural networks. However, the performance difference is less pronounced since models are retrained during validation months with a subset of samples.
Figure 9: F1-score over the test months for active learning on the (left) deduplicated (right) duplicated Drebin datasets
7.2 Variance in Results
Figure 9 presents the monthly F1 scores on the Drebin dataset for active learning settings, focusing on four models: SVM, XGBoost, MLP, and HCC, in both duplicated and deduplicated datasets. Separate hyperparameter tuning was conducted for SVM, XGBoost, and MLP, while for HCC, we used the artifacts provided in [15] for the duplicated setting, as it aligns with their study.
In the deduplicated dataset, XGBoost performs best in most test months. With an F1 score difference of at least 1% considered "significant," XGBoost outperforms HCC in 15 out of 18 test months, while HCC leads in 2 months, with one month showing less than a 1% difference. On the original duplicated dataset, HCC surpasses XGBoost in the global average F1 metric by 3.4%. However, a monthly analysis reveals a different picture: XGBoost outperforms HCC in 9 months, while HCC leads in 8 months. This highlights the importance of detailed monthly analyses in comparing malware detection models rather than relying solely on a single metric [35].
8 Number of Trials and Performance Convergence
We plot the performance on the validation set as the number of hyperparameter search trials increases across different datasets. This is shown for the offline learning setting in Figure 10 and for the active learning setting in Figure 11. Most models converge by 100 trials in the offline learning setting. In the active learning setting, convergence occurs even earlier, suggesting that initial hyperparameters have less impact on the results. While the performance trend on the validation set can be a good indicator of model performance, this is not always the case. For example, the RF model seems to perform much better on the validation set of the Drebin dataset for both offline and active learning settings than on the test set. Therefore, using a separate validation split for model selection in malware detection is crucial.
Figure 10: Performance on validation set as number of hyperparameter trials is increased for offline learning
Figure 11: Performance on validation set as number of hyperparameter trials is increased for active learning
9 Recommendations for Better Reproducibility
Based on our reproducibility study of Android malware detection, we propose a set of recommendations to enhance the reproducibility of reported results. While demonstrated in the context of Android malware detection, these guidelines apply to other security domains where temporal dataset evaluation is essential for fair assessment of model performance in real-world settings. These recommendations should be used alongside existing guidelines for machine learning in security [6, 41] and Android malware detection [42] to improve reproducibility in future work.
  • Datasets should be deduplicated to provide a more accurate estimation of model performance and improve reproducubility of research.
  • Datasets should be split into non-overlapping train, validation, and test sets for model selection and hyperparameter tuning. Once tuning is complete, the validation set can be merged with the training data in offline learning settings to estimate model performance on test data.
  • Proper hyperparameter tuning for baseline machine learning models is crucial to understanding improvements introduced by novel approaches. For Android malware detection, XGBoost is a strong benchmark for offline and active learning settings.
  • For neural networks, reporting performance using multiple random seeds on the test set is important to account for variance introduced by random weight initialization.
  • Instead of relying on a single metric, a thorough analysis should be performed on results obtained on a month-by-month basis to compare different models and identify any potential bias leading to flawed conclusion.
10 Limitations
In this section, we discuss the potential limitations of our work. We focus solely on exact duplicates in model input, ignoring cases of applications with similar functionalities that may result in near duplicates in the feature space [4]. Even with this strict definition, up to 50% duplicates can be present in publicly available Android malware detection datasets. Extending the duplicate analysis to other domains, such as Windows malware, might require near-duplicate analysis, such as Windows malware [5, 54], which use different features.
Our study focused on two datasets with different feature sets, but many other feature sets and datasets have been proposed in the literature [36]. Our goal was to highlight how duplicates and other methodological pitfalls can affect results rather than providing a thorough comparison of different data and feature sets. Our findings on model selection, robust baselines, and evaluation strategies to enhance reproducibility apply to other feature sets and domains and can be applied in the future.
Datasets with labels aggregated from sources such as VirusTotal may have issues with ground truth annotations, as noted in prior works [3], but this is beyond our study’s scope. We focus on offline and active learning for malware detection. In the active learning setting, we retrain the model monthly with new samples, as done in prior work. However, this may be inefficient if model performance has not degraded, and a drift detection model may be needed to determine when retraining is necessary [30]. Our analysis can be extended to datasets and models used for such tasks.
11 Conclusion
In this paper, we have critically examined the challenges of reproducibility and replicability in Android malware detection, focusing on key pitfalls in dataset curation, model selection, and evaluation strategies. Our extensive evaluation across multiple machine learning models in offline and active learning settings reveals that issues such as data duplication, inadequate hyperparameter tuning, and biased model selection can lead to significant performance disparities. These findings highlight the need for more rigorous approaches to ensure the reliability of reported results. Our results indicate that, with proper tuning, simpler models can often match or surpass the performance of more complex methods, challenging the trend of relying solely on sophisticated architectures. To support future research, we provide detailed recommendations to enhance reproducibility in Android malware detection using ML, many of which can also be applied to other security tasks. Our open-source code offers a standardized, extensible framework for benchmarking new models, facilitating fairer and more reliable comparisons across studies. By making our framework available, we aim to contribute to a more transparent and robust research environment, thereby enhancing the reproducibility of future work in this field.
Appendix A Implementation Details
A.1 Environment
We conducted all experiments using Python 3.11.6 on Red Hat Enterprise Linux (Version 9.4). Neural network training was performed using a single NVIDIA A100 GPU. Table 6 lists the Python pip packages used in the implementation:
Pip Package Version Utility
scikit-learn 1.5.0 Used to train RF, SVM
torch 2.1.0 Used to train neural networks (MLP, SCC, HCC)
xgboost 2.1.0 Used to train XGBoost
numpy 1.26.4 Used to process datasets
pytorch-metric-learning 2.6.0 Used for creating data sampler for HCC
yacs 0.1.8 Configuration management for experiments
Table 6: Python pip packages used in the implementation
A.2 Code Structure
The code is organized into multiple key modules to facilitate easy extension:
  1. 1.
    Dataset: We preprocess data into a single .npz file containing features, labels, timestamps, and duplication indicators. The intra_split_dupes field marks duplicate within the current timestamp, referring to earlier indices or set to -1 if unique. The cross_split_dupes field tracks duplicates across splits, indicating the relevant split and index or none if unique. This format supports active and offline learning deduplication. The MalwareDataset class handles dataset loading and includes the HalfSampler method used in HCC [15], as well as a class for working with Triplet Datasets.
  2. 2.
    Models: This module implements six models, all inheriting from a BaseModel class, which takes params for hyperparameters and cfg for experiment-specific settings (e.g., output classes, GPU usage). Each model includes methods for fitting (fit), prediction (predict, predict_proba), and active learning sample selection (sample_active_learning).
  3. 3.
    Tasks: We separate offline and active learning into distinct tasks within the tasks module, each with specific training and evaluation procedures. Tasks use a yacs configuration [23], provided as a YAML file, specifying the task type, model, and relevant settings. This separation allows easy extension for other malware analysis tasks, such as family classification or concept drift detection, with minimal code changes.
Appendix B Hyperparameter Search
B.1 Hyperparameter Search-Space
The detailed hyperparameter search space is shown in Table 7. Most hyperparameter names follow standard terminologies in the scikit-learn, XGBoost, or PyTorch APIs. Below, we describe those with different or new terminology:
  • XGBoost/MLP/SCC - class-weight: This parameter, "scale_pos_weight" in XGBoost, defaults to 1 if set to False. When True, it adjusts for class imbalance by weighting the positive class (malware) according to the benign-to-malware ratio in the training data. For MLP and SCC, samples are either balanced in each batch with equal benign and malware samples (balance=True) or drawn randomly (balance=False).
  • MLP/SCC/HCC - cont_learning_epochs: In active learning with neural networks, initial and retraining epochs differ, similar to [15]. For HCC, it specifies the number of epochs; for MLP and SCC, it represents the fraction of the original training epochs.
  • HCC - cont_learning_lr: For HCC, the initial learning rate is varies the learning rate for each update, while MLP and SCC continue with the current rate using the Adam optimizer.
  • SCC/HCC - xent_lambda: This controls the binary cross-entropy loss weight in contrastive learning. We use a value of 100, as suggested in [15].
  • SCC/HCC - margin: This sets the margin for computing contrastive loss. We use a value of 10, following [15].
We adopt strategies from prior works on Android malware detection [15], machine learning on tabular data [39], and deep neural architectures [9] to design the hyperparameter search space.
While we mostly follow the hyperparameters for HCC from [15], there are a few minor differences. These choices ensure consistency with other neural network architectures and ease of implementation. We explore two architectures for the encoder and MLP layers, each instead of the single one used in their work. We include dropout as a hyperparameter instead of setting it to a fixed value of 0.2 in their study. Only the Adam optimizer is used during the active learning retraining phase. We employ a step learning rate scheduler and omit the cosine annealing scheduler. Regardless, we verify that our search space includes the best hyperparameter set reported in [15]; for instance, cosine annealing did not outperform step-based learning rates in any dataset in their results.
Model Hyperparameter Candidate Values
Random Forest n_estimators $2^{x}$, where $x\in\text{U}[5,10]$
max_depth $2^{y}$, where $y\in\text{U}[5,10]$
criterion {gini, entropy, log_loss}
class_weight {None, "balanced"}
SVM C $10^{z}$, where $z\in\text{U}[-4,3]$
class_weight {None, "balanced"}
XGBoost max_depth $2^{w}$, where $w\in\text{U}[3,7]$
alpha $10^{a}$, where $a\in\text{U}[-8,0]$
lambda $10^{b}$, where $b\in\text{U}[-8,0]$
eta $3.0\times 10^{c}$, where $c\in\text{U}[-2,-1]$
balance {True, False}
num_boost_round {100, 150, 200, 300, 400}
MLP mlp_layers {[100, 100], [512, 256, 128], [512, 384, 256, 128], [512, 384, 256, 128, 64]}
learning_rate $10^{d}$, where $d\in\text{U}[-5,-3]$
dropout $x$, where $x\in\text{U}[0.0,0.5]$
batch_size $2^{e}$, where $e\in$5, 6, 7, 8, 9, 10}
epochs {25, 30, 35, 40, 50, 60, 80, 100, 150}
optimizer {Adam}
balance {True, False}
cont_learning_epochs* {0.1, 0.2, 0.3, 0.4, 0.5}
SCC encoder_layers {[512, 256, 128], [512, 384, 256, 128]}
mlp_layers {[100], [100, 100]}
learning_rate $10^{f}$, where $f\in\text{U}[-5,-3]$
dropout $y$, where $y\in\text{U}[0.0,0.25]$
batch_size $2^{g}$, where $g\in$9, 10, 11}
epochs {25, 30, 35, 40, 50, 60, 80, 100}
xent_lambda {100}
margin {10}
optimizer {Adam}
balance {True, False}
cont_learning_epochs* {0.1, 0.2, 0.3, 0.4, 0.5}
HCC encoder_layers {[512, 256, 128], [512, 384, 256, 128]}
mlp_layers {[100], [100, 100]}
learning_rate {0.001, 0.003, 0.005, 0.007}
dropout $z$, where $z\in\text{U}[0.0,0.25]$
batch_size $2^{10}$
epochs {100, 150, 200, 250}
xent_lambda {100}
margin {10}
optimizer {Adam, SGD}
scheduler_step {10}
scheduler_gamma {0.5, 0.95}
cont_learning_lr* {0.01, 0.05}
cont_learning_epochs* {50, 100}
Table 7: Hyperparameter Search Spaces for Different Models. U indicates sampling from a uniform random distribution over values in the range. * indicates hyperparameters that are only used during active learning retraining phase.
B.2 Influence of Individual Hyperparameters
Figure 12 illustrates the impact of individual hyperparameters on the validation set performance for the XGBoost model in the offline learning setting.
Figure 12: Influence of individual hyperparameters on different datasets for XGBoost